# Fab-ME: A Vision State-Space and Attention-Enhanced Framework for Fabric Defect Detection

### v1

@@ Abstract
Effective defect detection is critical for ensuring the quality, functionality, and economic value of textile products. However, existing methods face challenges in achieving high accuracy, real-time performance, and efficient global information extraction. To address these issues, we propose Fab-ME, an advanced framework based on YOLOv8s, specifically designed for the accurate detection of 20 fabric defect types. Our contributions include the introduction of the cross-stage partial bottleneck with two convolutions (C2F) vision state-space (C2F-VMamba) module, which integrates visual state-space (VSS) blocks into the YOLOv8s feature fusion network neck, enhancing the capture of intricate details and global context while maintaining high processing speeds. Additionally, we incorporate an enhanced multi-scale channel attention (EMCA) module into the final layer of the feature extraction network, significantly improving sensitivity to small targets. Experimental results on the Tianchi fabric defect detection dataset demonstrate that Fab-ME achieves a 3.3% improvement in mAP@0.5 compared to the original YOLOv8s, validating its effectiveness for precise and efficient fabric defect detection.
@@ I Introduction
In this paper, we present Fab-ME, an enhanced framework based on YOLOv8s, for detecting 20 types of fabric defects. To address challenges in detecting large targets, we integrate visual state-space (VSS) blocks [25] into the cross-stage partial bottleneck with two convolutions (C2F) module, forming the C2F-VMamba module. This integration improves global context understanding and detail preservation while maintaining real-time performance. Additionally, an enhanced multi-scale channel attention (EMCA) module is incorporated into the feature extraction network, boosting sensitivity to small targets. Experimental results on the Tianchi fabric defect detection dataset show a 3.3% improvement in mAP@0.5 over YOLOv8s, with real-time processing capabilities. The main contributions of this work are summarized as follows:
@@
  • The Fab-ME framework achieves a 3.3% mAP@0.5 improvement on the Tianchi dataset while maintaining real-time performance.
@@ IV Conclusion
We proposed Fab-ME, a real-time fabric defect detection framework based on YOLOv8s, designed to improve accuracy, efficiency, and global feature extraction. Fab-ME integrates the C2F-VMamba module, leveraging VSS blocks for enhanced global context and detail capture, and the EMCA module, which improves multi-scale sensitivity and small target detection. Experiments on the Tianchi dataset showed a 3.3% mAP@0.5 improvement over YOLOv8s, demonstrating the effectiveness of our approach. Future work will address challenges with subtle defects and explore advanced attention mechanisms and diverse datasets to further enhance robustness and detection performance.

### v2

@@ Abstract
Effective defect detection is critical for ensuring the quality, functionality, and economic value of textile products. However, existing methods face challenges in achieving high accuracy, real-time performance, and efficient global information extraction. To address these issues, we propose Fab-ME, an advanced framework based on YOLOv8s, specifically designed for the accurate detection of 20 fabric defect types. Our contributions include the introduction of the cross-stage partial bottleneck with two convolutions (C2F) vision state-space (C2F-VMamba) module, which integrates visual state-space (VSS) blocks into the YOLOv8s feature fusion network neck, enhancing the capture of intricate details and global context while maintaining high processing speeds. Additionally, we incorporate an enhanced multi-scale channel attention (EMCA) module into the final layer of the feature extraction network, significantly improving sensitivity to small targets. Experimental results on the Tianchi fabric defect detection dataset demonstrate that Fab-ME achieves a 3.5% improvement in mAP@0.5 compared to the original YOLOv8s, validating its effectiveness for precise and efficient fabric defect detection.
@@ I Introduction
In this paper, we present Fab-ME, an enhanced framework based on YOLOv8s, for detecting 20 types of fabric defects. To address challenges in detecting large targets, we integrate visual state-space (VSS) blocks [25] into the cross-stage partial bottleneck with two convolutions (C2F) module, forming the C2F-VMamba module. This integration improves global context understanding and detail preservation while maintaining real-time performance. Additionally, an enhanced multi-scale channel attention (EMCA) module is incorporated into the feature extraction network, boosting sensitivity to small targets. Experimental results on the Tianchi fabric defect detection dataset show a 3.5% improvement in mAP@0.5 over YOLOv8s, with real-time processing capabilities. The main contributions of this work are summarized as follows:
@@
  • The Fab-ME framework achieves a 3.5% mAP@0.5 improvement on the Tianchi dataset while maintaining real-time performance.
@@ IV Conclusion
We proposed Fab-ME, a real-time fabric defect detection framework based on YOLOv8s, designed to improve accuracy, efficiency, and global feature extraction. Fab-ME integrates the C2F-VMamba module, leveraging VSS blocks for enhanced global context and detail capture, and the EMCA module, which improves multi-scale sensitivity and small target detection. Experiments on the Tianchi dataset showed a 3.5% mAP@0.5 improvement over YOLOv8s, demonstrating the effectiveness of our approach. Future work will address challenges with subtle defects and explore advanced attention mechanisms and diverse datasets to further enhance robustness and detection performance.